On April 1, 2021, Austere Corporation issued $350,000 of 10% bonds at 107. Each $1,000 bond was sold with 40 detachable stock warrants, each permitting the investor to purchase one share of common stock for $17. On that date, the market value of the common stock was $15 per share and the market value of each warrant was $2. Austere should record what amount of the proceeds from the bond issue as an increase in liabilities?

Multiple Choice

  • $271,000.

  • $0.

  • $346,500.

  • $471,800.
Answer #1

Solution:

Proceed from issue of bond = $350,000*107% = $374,500

Fair value of warrant issued = 350*40*$2 = $28,000

Austere should record amount of the proceeds from the bond issue as an increase in liabilities = $374,500 - $28,000 = $346,500

Hence 3rd option is correct.